Circular No. 8969

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 R1 (GARRADD) G. J. Garradd reports the discovery of a comet with a 15′′ coma (more

prominent on the northern side) on CCD images taken with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, G. Sostero, Remanzacco, Italy, reports that P. Camilleri, E. Guido, and he co-added 24 unfiltered 120-s CCD exposures, obtained remotely with a 0.35-m f/6.7 reflector of Grove Creek-Skylive Observatory at Trunkey, NSW, Australia, on Sept. 2.71–2.79 UT, showing a coma ∼ 8′′ in diameter, elongated toward p.a. ∼ 30^{o}. C. Jacques and E. Pimentel, Belo Horizonte, Minas Gerais, Brazil, obtained CCD images on Sept. 4.1 with a 0.30-m f/3.0 reflector that show a coma diameter of 25′′.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.2.49906  21^{h}41^{m}12.^{s}73  −47^{o}47′56′′.4  18.5

The available astrometry, very preliminary parabolic orbital elements (T = 2008 Nov. 8.969 TT, q = 1.70548 AU, ω = 336^{o}.687, Ω = 27^{o}.273, i = 19^{o}.027, equinox 2000.0), and an ephemeris appear on MPEC 2008-R22.

V1647 ORIONIS C. Aspin, Institute for Astronomy, University of Hawaii (UH), reports

that an optical spectrum of V1647 Ori was obtained on Aug. 30 UT using the UH 2.2-m telescope (+ SNIFS integral-field spectrograph). Using the Spitzer flux-to-magnitude converter (http://ssc.spitzer.caltech.edu/ tools/magtojy/) and the flux at 650.0 nm from the calibrated SNIFS spectrum (2.1 × 10^{−16} erg s^{−1} cm^{−2} Å^{−1}) suggests that, on this date, V1647 Ori had a Johnson R magnitude of 17.3. This is around 6 magnitudes brighter than the measured quiescent brightness of r′ = 23.3. The spectrum has a spectral resolution of around R = 1300 and shows a very red continuum with strong Hα emission. The Hα emission line exhibited a P-Cyg profile with a blue-shifted absorption trough extending to around 500 km/s from the line rest wavelength. In addition, the red Ca II triplet lines were strongly in emission. McNeil’s nebula (cf. IAUC 8284, 8354), illuminated by V1647 Ori, was also strongly visible. This confirms that V1647 Ori has undergone another strong outburst (cf. IAUC 8968) similar to the one observed in 2003, the previous event lasting around 26 months.

2008 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green